# DPOD: Domain-Specific Prompt Tuning for Multimodal Fake News Detection

### v3

@@ Abstract
The spread of fake news using out-of-context images has become widespread and is a relevant problem in this era of information overload. Such out-of-context fake news may arise across different domains like politics, sports, entertainment, etc. In practical scenarios, an inherent problem of imbalance exists among news articles from such widely varying domains, resulting in a few domains with abundant data, while the rest containing very limited data. Under such circumstances, it is imperative to develop methods which can work in such varying amounts of data setting. In this work, we explore whether out-of-domain data can help to improve out-of-context misinformation detection (termed here as multi-modal fake news detection) of a desired domain, to address this challenging problem. Towards this goal, we propose a novel framework termed DPOD (Domain-specific Prompt-tuning using Out-of-Domain data). First, to compute generalizable features, we modify the Vision-Language Model, CLIP to extract features that helps to align the representations of the images and corresponding text captions of both the in-domain and out-of-domain data in a label-aware manner. Further, we propose a domain-specific prompt learning technique which leverages the training samples of all the available domains based on the extent they can be useful to the desired domain. Extensive experiments on a large-scale benchmark dataset, namely NewsCLIPpings demonstrate that the proposed framework achieves state of-the-art performance, significantly surpassing the existing approaches for this challenging task. Code will be released on acceptance.
Keywords:
Multi-modal fake news detection Prompt tuning Vision-language models
@@ 1 Introduction
In this digital world, social media has become the main source of information for an increasing fraction of the population. Thus, it is very important that the news reaching the masses is authentic, since fake news can have serious consequences, like manipulating public opinion, stirring up conflicts, and even impacting financial markets. With the internet connecting billions of people worldwide, fake news can spread at an alarming speed, thus it is equally important to identify it quickly and efficiently. One increasingly common form of fake news is the out-of-context use of images, where a real image is paired with a false caption before dissemination. It is very difficult to detect such fake news, since the image is real, i.e., not manipulated or generated, and it is only recently that researchers have started to address this challenging, but very important problem [2][1].
In a real-world setting, there is an implicit skewness present in data across varying domains. For example, there is an increased dominance of news domains like politics over certain specific topics like education. However, one maybe interested in news from a particular domain, e.g., a doctor may be more interested in healthcare news, whereas a government official may be more interested in political news. Due to the presence of such widely varying data quantities across different domains, it is difficult to correctly predict the veracity of a news sample belonging to a particular domain of interest. However, news articles may often share overlapping traits, with different domains intersecting with each other. Hence, for a target domain of choice, we can leverage information from other domains to address this problem of imbalance.
@@
In this work, we propose a novel framework termed DPOD (Domain-specific Prompt-tuning using Out-of-Domain data) for out-of-context misinformation detection by leveraging both image and caption pairs, which we term here as MFND or Multi-modal Fake News Detection. It is evident that training a model on the samples of a specific domain will yield good performance when evaluated on the same domain. However, this is not practical, since in addition to training different models for different domains, training on domains with less data may be difficult. Towards this goal, we develop a generalized framework, where we consider data from all domains while training. The model is trained such that it is simultaneously tailored to each individual domain, by leveraging information from all the available domains. As an example, for MFND on a particular domain of interest like politics, we effectively utilize out-of-domain data from healthcare, international relations, etc. to improve its performance. The DPOD framework first utilizes label-aware alignment of the available image-text pairs of all the domains to learn generalizable features, irrespective of the domain of interest. In our work, we harness the capabilities of a large Vision Language Model (CLIP) [18] to comprehend interwoven text and image features. Next, a semantic domain vector is computed for each of the training data domains, which incorporates the similarity of a particular domain with all the other available training domains. Further, this is utilized to learn a domain-specific prompt, along with a set of generic learnable prompts, which is finally used for predicting whether a given image-caption pair is real or fake. During inference, the model specialized to the test data can be utilized, even when the domain label of the test data is unknown. Extensive experiments on a large-scale benchmark dataset, namely NewsCLIPpings [15] demonstrate the effectiveness of DPOD over state-of-the-art approaches. The contributions of this work can be summarized as follows:
  1. 1.
@@
    We propose a novel framework, termed DPOD, which effectively utilizes out-of-domain data for improving the performance of in-domain data, hence addressing the issue of varying amounts of data across different domains.
  2. 2.
@@
    We utilize a label-aware loss to learn a generalizable model, and propose a novel framework to learn domain-specific prompts using out-of-domain data.
  3. 3.
@@
    This helps in quick model deployment for the target domain, even when it contains limited domain-specific annotated data, thus being effective for scarcely annotated data scenarios.
  4. 4.
@@
    Extensive experiments show that the proposed DPOD outperforms the existing approaches in different settings, thus achieving the new state-of-the-art.
@@ 2 Related Work
Multi-modal Fake News Detection: Recently due to the rapid spread of fake news using out-of-context real images and texts/captions, researchers have started to address this socially relevant important problem. [15] proposed a novel dataset named NewsCLIPpings, derived from the Visual News dataset [14] and suggested a baseline solution using CLIP and VisualBERT feature extractor. The SAFE framework [29] utilises Text-CNN and VGG [20] model for extracting features from the texts and their associated images which are then used to predict the fake news. Similarly, [22] use BERT [5] as text encoder to extract textual features and use VGG-19 as image encoder to extract visual features and then finally fuse them together for classification. Most of these existing methods fail to perform well on unseen events. [23] trained an event discriminator network concurrently with multi-modal fake news detector to eliminate event-specific characteristics while retaining the common features among different events. The work in [19] jointly preserves domain-specific and cross-domain knowledge in news records to detect fake news from different domains by using combination of domain-specific and cross-domain features. In [24], meta-learning and neural process methods are integrated to achieve high performance even on events with limited labeled data.
@@
The emergence of large scale vision-language models (VLMs) which learns information from both images and texts, has proved to be immensely useful for computer vision tasks. Models like CLIP [18] and ALIGN [11] have been trained in a contrastive manner on large scale datasets, such that they learn rich representations between images and texts. [30] used CLIP along with image encoder (ResNet) and text encoder (BERT) to extract the multimodal features and used attention over different modalities to get the final features for classification. Recently, [17] proposed a novel DT-Transformer framework for MFND using CLIP encoders and a Transformer-based detector instead of using a binary classifier. Some recent works also use social media related information to detect the fake news [19][1][25]. [19] proposed a framework which uses image, text and propagation graph for MFND and [25] suggested a graph model based on CLIP embeddings. A Consistency-Checking Network is proposed in [1] to measure the consistency of the image and text pairs along with the evidences collected from the internet.
@@
Self-Supervised Learning: Self-supervised learning leverages the inherent structure of data to generate its own supervisory signals without relying on large labeled datasets. [4] have showed that the combination of data augmentations has a crucial impact on useful predictive tasks. Supervised contrastive loss [13] leverages label information for self-supervised representation learning. It pulls together clusters of points belonging to the same class while pushing apart clusters of samples from different classes. Recently, self-supervision has been used successfully for out-of-context misinformation detection using images and text. In [2], self-supervised learning is used to align individual objects in an image with the corresponding textual claims, where, given two captions, the goal is to detect whether an image is being used out-of-context. The state-of-the-art method  [16] propose a Self-Supervised Distilled Learner, which uses self-supervised learning to detect fake news, where they propose a Teacher network to guide a Student network in mimicking similar decision patterns. Though we also utilize self-supervision, our proposed DPOD framework is much simpler, yet more effective, and does not rely on teacher-student model.
@@
Vision-Language Models and Prompt Tuning: Since, adapting large VLM’s like CLIP to downstream tasks may prove to be difficult, methods like prompt learning have come up as a popular research direction. The text instructions given to the text encoder, are called “Prompts”. Prompt learning is an alternative to handcrafted prompts, where the prompt tokens are learned during the finetuning stage. Recently, many works have proposed prompt learning techniques such as CoOp [28], CoCoOp [27], Maple [12], etc., which increases the generalizability of the VLMs. Some works specifically address the domain adaptation problem, for example, [26] proposes a domain-invariant prompt tuning, which can generalize to unseen domains. [9] uses prompt tuning in a unimodal context (in language models), whereas, [3] uses only visual prompt tuning. The works in [8, 21, 6] learns prompts specifically for each domain. To the best of our knowledge, ours is the first work to explore prompt-tuning for the challenging MFND task.
@@ 3 Problem Definition and Model Overview
The goal of the proposed framework is to leverage out-of-domain training samples when testing on a certain target domain. Let the train dataset be denoted as $\mathcal{S}_{train}=\{{D}_{1},{D}_{2},...,{D}_{n}\}$. ${D}_{i}$ denotes each domain and can be written as, ${D}_{i}=\{({I}_{k},{T}_{k},{y}_{k})\},k=1,2,...,{n}_{{D}_{i}}$, where ${n}_{{D}_{i}}$ is the number of samples in domain ${D}_{i}$. ${I}_{k}$ and ${T}_{k}$ denote the $k^{th}$ image and text pair, and ${y}_{k}$ denotes the ground truth, i.e., fake or true news. The news samples can originate from various sources (or domains) such as politics, healthcare, sports, etc. We assume that we know the domain of the training examples, though we show empirically that the proposed framework can inherently handle domain label inconsistencies that is expected from crowd-sourced data annotation. We assume that an user is interested in a specific domain $D\in\{{D}_{1},{D}_{2},...,{D}_{n}\}$ during testing, though our model training happens only once irrespective of the desired target domain.
During testing, we want to predict whether a given image-text pair is fake or real. First, we briefly describe the base model used.
@@
Stage 1: Label-Aware Alignment of Multi-modal data: In this stage, we train all the layers of the original CLIP model in an end-to-end manner using the label-aware alignment loss on all the training data, irrespective of their domains, to obtain the aligned clip, termed here as A-CLIP.
@@
Stage 3: Domain-Specific Prompt tuning with Out-of-Domain data: With the above settings, we next propose a prompting strategy during training. Following previous works [28][27], we append a prefix prompt to the text caption. However, a fixed prompt like “A photo of” [18], may not be the best prefix for all the captions. Hence, we make the prefix prompt tokens learnable as $\{{V}_{1},{V}_{2},{V}_{3}\}$. Additionally, we pass the semantic domain vector though a linear layer $\mathcal{F}_{\theta}$, and get a domain-specific prompt token, which is appended to the prefix prompt tokens. The complete text captions appended with the learnable prompts, and the corresponding images are passed though the frozen encoders to create the final joint embedding, which is subsequently classified into fake or real news.
@@ 4.1 Stage 1: Label-Aware Alignment
Since our goal is to predict the veracity of news (image-caption pair) in a particular domain ($D$, say politics), the ideal scenario is when vast amounts of annotated data for domain $D$ are available. In real scenarios, getting large amounts of annotated data from each domain is challenging. Our hypothesis is that there might be overlapping concepts/objects in the data from different domains, like healthcare, entertainment, which can be utilized to mitigate the adverse effect of less amount of training data of the desired domain.
Using self-supervision as a warm-up stage for model training has been used successfully in previous works like [2, 16]. In contrast to these approaches, where both the training and test sets contain data from all (but same) domains, we explore whether self-supervision from out-of-domain data can benefit the performance of in-domain data for MFND task. Here, the original CLIP model is trained to learn relations between texts and images from all the domains, but in a label-aware manner. Specifically, we use a multi-modal contrastive loss, such that the backbone model learns to bring closer the augmentations of the image and the corresponding text when the news is real and can separate them when the news is fake.
For label-aware alignment, initially, four augmentations, namely, color jitter, random crop, horizontal flipping and normalisation are generated per image. Then the CLIP embeddings of the original image, text and the image augmentations are generated. Suppose an image and the corresponding text be denoted as $I_{k}$ and $T_{k}$ respectively and the image augmentations be denoted by $I_{ak},a=1,\ldots,4$. We denote the corresponding embeddings as $\{\hat{I}_{k},\hat{T}_{k},\hat{I}_{ak}\}$. For the consistent (real) image-text pairs, we want to bring the embeddings of the image ($\hat{I}_{k}$), its augmentations ($\hat{I}_{ak},a=1,\ldots,4$) and the corresponding text ($\hat{T}_{k}$) close to one another. Thus for the positive pair, we have two combinations, namely image-image pair and image-text pair. The corresponding label-aware alignment losses for the $i^{th}$ image augmentation of $k^{th}$ input sample are calculated as:
@@
$\mathcal{D}_{\text{true}}=\sum_{\begin{subarray}{c}t=1\\ t\neq k\end{subarray}}^{N}\left(\sum_{a=1}^{4}\exp(\text{sim}(\hat{I}_{ik},\hat{I}_{at}))+\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{t}))+\exp(\text{sim}(\hat{T}_{k},\hat{T}_{t}))\right)$
@@
$\mathcal{D}_{\text{fake}}=\sum_{\begin{subarray}{c}t=1\\ t\neq k\end{subarray}}^{N}\left(\sum_{a=1}^{4}\exp(\text{sim}(\hat{I}_{ik},\hat{I}_{at}))+\exp(\text{sim}(\hat{T}_{k},\hat{T}_{t}))\right.$
$+\left.\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{t}))\right)+\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{k}))$

### v4

# Can Out-of-Domain data help to Learn Domain-Specific Prompts for Multimodal Misinformation Detection?

@@ Abstract
Spread of fake news using out-of-context images and captions has become widespread in this era of information overload. Since fake news can belong to different domains like politics, sports, etc. with their unique characteristics, inference on a test image-caption pair is contingent on how well the model has been trained on similar data. Since training individual models for each domain is not practical, we propose a novel framework termed DPOD (Domain-specific Prompt tuning using Out-of-domain data), which can exploit out-of-domain data during training to improve fake news detection of all desired domains simultaneously. First, to compute generalizable features, we modify the Vision-Language Model, CLIP to extract features that helps to align the representations of the images and corresponding captions of both the in-domain and out-of-domain data in a label-aware manner. Further, we propose a domain-specific prompt learning technique which leverages training samples of all the available domains based on the extent they can be useful to the desired domain. Extensive experiments on the large-scale NewsCLIPpings and VERITE benchmarks demonstrate that DPOD achieves state of-the-art performance for this challenging task. Code: https://github.com/scviab/DPOD.
1BBI Inc., 2Indian Institute of Science, Bangalore
amartya.bhattacharya@bbiinc.ai,
{debarshib, surajnagaje, anmolasati, vikasverma, somabiswas}@iisc.ac.in
**footnotetext: Equal contribution.$\dagger$\$$\dagger$\$footnotetext: Work done at the Indian Institute of Science, Bangalore.
@@ 1 Introduction
In this digital world, social media has become the main source of information for an increasing fraction of the population. Thus, it is very important that the news reaching the masses is authentic, since fake news can have serious consequences, like manipulating public opinion, stirring up conflicts, and even impacting financial markets. One increasingly common form of fake news is the out-of-context use of images, where a real image is paired with a false caption before dissemination. It is very difficult to detect such fake news, since the image is real, i.e., not manipulated or generated, and it is only recently that researchers have started to address this challenging problem [2][1].
In real-world, fake news can belong to any domain like healthcare, politics, entertainment, etc. Obviously, the accuracy of authentication of a test image-text pair belonging to a particular domain depends on how well that domain is represented in the training data. Training different models for different domains is not a practical solution, since it assumes abundant availability of data from every domain, in addition to not effectively utilizing the inter-relations and overlapping traits that may be present among them.
@@
Towards this goal, we propose a novel, generalized framework termed DPOD (Domain-specific Prompt-tuning using Out-of-Domain data) for out-of-context misinformation detection, also termed as Multi-modal Fake News Detection (MFND). The model is trained using data from all the available domains, such that it can then be simultaneously tailored to each individual domain, by leveraging information from all the other domains. As an example, for MFND on a particular domain of interest like politics, we effectively utilize out-of-domain data from healthcare, international relations, etc. to improve its performance. The DPOD framework first utilizes label-aware alignment of the available image-text pairs of all the domains to learn generalizable features, irrespective of the domain of interest. Here, we harness the capabilities of a large Vision Language Model (CLIP) [20] to comprehend interwoven text and image features. Next, a semantic domain vector is computed for each of the training data domains, which incorporates the similarity of a particular domain with all the other available training domains. Further, this is utilized to learn a domain-specific prompt, along with a set of generic learnable prompts, which is finally used for predicting whether a given image-caption pair is real or fake. During inference, the model specialized to the test domain can be utilized, even when the test domain is unknown. Extensive experiments on two benchmark datasets, namely NewsCLIPpings [15] and VERITE [18], demonstrate the effectiveness of DPOD over state-of-the-art approaches. The contributions of this work can be summarized as follows:
@@
1) We propose a novel framework, termed DPOD, which effectively utilizes out-of-domain data for improving the performance of in-domain data.
@@
2) We utilize a label-aware loss to learn a generalizable model, and propose a novel framework to learn domain-specific prompts using out-of-domain data.
@@
3) This helps in quick model deployment for the target domain, even with limited domain-specific annotated data, thus being effective for scarcely annotated data scenarios.
@@
4) Extensive experiments show that the proposed DPOD outperforms the existing approaches in different settings, thus achieving the new state-of-the-art.
@@ 2 Related Work
Multi-modal Fake News Detection: To address the MFND task, [15] proposed NewsCLIPpings dataset, derived from VisualNews [14], and suggested using CLIP and VisualBERT feature extractor. SAFE [32] utilizes Text-CNN and VGG [22] model for extracting features from the texts and images which are then used to predict the fake news. Similarly, [24] use BERT [5] and VGG-19 to extract textual and visual features and then fuse them for classification. Most of these methods fail to perform well on unseen events. [25] trained an event discriminator network concurrently with multi-modal fake news detector to eliminate event-specific characteristics while retaining common features among different events.  [21] jointly preserves domain-specific and cross-domain knowledge in news records to detect fake news from different domains. In [26], meta-learning and neural process methods are integrated to achieve high performance even on events with limited labeled data.
@@
Recently, large scale vision-language models (VLMs) like CLIP [20] and ALIGN [11], which learns information from both images and texts, have greatly advanced computer vision tasks. [33] used CLIP along with ResNet and BERT to extract multimodal features and used attention over different modalities to get the final features for classification. [17] proposed a novel framework for MFND using CLIP encoders and a Transformer-based detector. Some recent works also use social media related information to detect fake news [21][1][28]. [21] uses image, text and propagation graph for MFND and [28] suggested a graph model based on CLIP embeddings.  [1] proposes a Consistency-Checking Network to measure the consistency of the image-text pairs along with evidences collected from the internet. [27] uses a combination of fine-grained and course-grained attention network along with deep reinforcement learning, while very recently [19] has explored the use of MLLMs in this problem. In our work, we do not assume the availability of open-domain information.
@@
Self-Supervised Learning (SSL): SSL leverages the inherent structure of data to generate its own supervisory signals without relying on large labeled datasets. [4] have showed that the combination of data augmentations has a crucial impact on predictive tasks. Supervised contrastive loss [13] leverages label information for self-supervised representation learning. It pulls together clusters of points belonging to the same class while pushing apart clusters of samples from different classes. Recently, self-supervision has been used successfully for out-of-context misinformation detection using images and text. In [2], SSL is used to align individual objects in an image with the corresponding textual claims, where, given two captions, the goal is to detect whether an image is being used out-of-context. The state-of-the-art method  [16] propose a Self-Supervised Distilled Learner, which uses self-supervised learning to detect fake news, where they propose a Teacher network to guide a Student network in mimicking similar decision patterns. Though we also utilize self-supervision, our proposed DPOD framework is much simpler, yet more effective, and does not rely on teacher-student model.
@@
Vision-Language Models and Prompt Tuning: Adapting large VLMs like CLIP to downstream tasks is challenging. Parameter efficient finetuning techniques like prompt learning is a popular approach, where the text prompt tokens are learned during the finetuning stage. Recently, many works have proposed prompt learning techniques such as CoOp [31], CoCoOp [30], Maple [12], etc., which increases the generalizability of the VLMs. Some works specifically address the domain adaptation problem, e.g., [29] proposes domain-invariant prompt tuning, which can generalize to unseen domains. [9] uses unimodal prompt tuning (in language models), whereas, [3] uses only visual prompt tuning. The works in [8, 23, 6] learns prompts specifically for each domain. In this work, we explore how prompt learning can be used for the challenging MFND problem.
@@ 3 Problem Definition and Model Overview
The goal is to develop a generalized MFND model, which effectively incorporates the domain information of the training data and relation between the different domains during training. Let the train dataset be denoted as $\mathcal{S}_{train}=\{{D}_{1},{D}_{2},...,{D}_{n}\}$. ${D}_{i}$ denotes each domain and can be written as, ${D}_{i}=\{({I}_{k},{T}_{k},{y}_{k})\},k=1,2,...,{n}_{{D}_{i}}$, where ${n}_{{D}_{i}}$ is the number of samples in domain ${D}_{i}$. ${I}_{k}$ and ${T}_{k}$ denote the $k^{th}$ image and text pair, and ${y}_{k}$ denotes the ground truth, i.e., fake or true news. The news samples can originate from various sources (or domains) such as politics, healthcare, sports, etc. We assume that we know the domain of the training examples, though we show empirically that the proposed framework can inherently handle domain label inconsistencies that is expected from crowd-sourced data annotation. The model is trained once using all the input domains, such that depending on the domain of the test image-text pair $D\in\{{D}_{1},{D}_{2},...,{D}_{n}\}$, the domain specific model is utilized for best performance. First, we briefly describe the base model used.
@@
Stage 1: Label-Aware Alignment of Multi-modal data: In this stage, we train the encoders of the original CLIP model in an end-to-end manner using the label-aware alignment loss on the training data from all domains, to obtain the aligned clip, termed here as A-CLIP.
@@
Stage 3: Domain-Specific Prompt tuning: With the above settings, we next propose a prompting strategy during training. Following previous works [31][30], we append a prefix prompt to the text caption. However, a fixed prompt like “A photo of” [20], may not be the best prefix for all the captions. Hence, we make the prefix prompt tokens learnable as $\{{V}_{1},{V}_{2},{V}_{3}\}$. Additionally, we pass the semantic domain vector though a linear layer $\mathcal{F}_{\theta}$, and get a domain-specific prompt token, which is appended to the prefix prompt tokens. The complete text captions with the learnable prompts, and the corresponding images are passed though the frozen encoders to create the final joint embedding, which is subsequently classified into fake or real news.
Figure 2: Computation of the joint embedding of the images and text for a particular domain. These embedding are used in Stage 2 and 3 of our DPOD framework.
@@ 4.1 Stage 1: Label-Aware Alignment
If we want to predict the veracity of news (image-caption pair) in a particular domain ($D$, say politics), the ideal scenario is when vast amounts of annotated data for domain $D$ are available. In real scenarios, getting large amounts of annotated data from each domain is challenging. But, there might be overlapping concepts or objects in the data from related domains, like politics, international news, which can be utilized to bolster the learning for each domain.
Using self-supervision as a warm-up stage for model training has been used successfully in previous works like [2, 16]. In contrast to these approaches, here, the original CLIP model is trained to learn relations between texts and images from all the domains, but in a label-aware manner. Specifically, we use a multi-modal contrastive loss, such that the backbone model learns to bring closer the augmentations of the image and the corresponding text when the news is real and can separate them when the news is fake.
For label-aware alignment, initially, four augmentations, namely, color jitter, random crop, horizontal flipping and normalization are generated per image. Then the CLIP embeddings of the original image, text and the image augmentations are generated. Suppose an image and the corresponding text be denoted as $I_{k}$ and $T_{k}$ respectively and the image augmentations be denoted by $I_{ak},a=1,\ldots,4$. We denote the corresponding embeddings as $\{\hat{I}_{k},\hat{T}_{k},\hat{I}_{ak}\}$. For the consistent (real) image-text pairs, we want to bring the embeddings of the image ($\hat{I}_{k}$), its augmentations ($\hat{I}_{ak},a=1,\ldots,4$) and the corresponding text ($\hat{T}_{k}$) close to one another. Thus for the positive pair, we have two combinations, namely image-image pair and image-text pair. The corresponding label-aware alignment losses for the $i^{th}$ image augmentation of $k^{th}$ input sample are calculated as:
@@
$\mathcal{D}_{true}=\sum_{\begin{subarray}{c}t=1\\ t\neq k\end{subarray}}^{N}\biggl{(}\sum_{a=1}^{4}\exp(\text{sim}(\hat{I}_{ik},\hat{I}_{at}))+\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{t}))+$
$\exp(\text{sim}(\hat{T}_{k},\hat{T}_{t}))\biggl{)}$
@@
$\mathcal{D}_{fake}=\sum_{\begin{subarray}{c}t=1\\ t\neq k\end{subarray}}^{N}\biggl{(}\sum_{a=1}^{4}\exp(\text{sim}(\hat{I}_{ik},\hat{I}_{at}))+\exp(\text{sim}(\hat{T}_{k},\hat{T}_{t})+$
$\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{t}))\biggl{)}+\exp(\text{sim}(\hat{I}_{ik},\hat{T}_{k}))$